OpenStudy (anonymous):

Evaluate r pie 2 for for = 3.14 and r = 3. help me :(

OpenStudy (anonymous):

just mutiply 2 3.14 and 3

OpenStudy (anonymous):

Do you mean \[\pi r ^2\]

OpenStudy (anonymous):

If so, then \[\pi r^2= 3.14 \times r \times r = 3.14 \times 3 \times 3\]